# MemANNS: Enhancing Billion-Scale ANNS Efficiency with Practical PIM Hardware

### v1

@@ 4.4. Top-K Pruning
The above optimizations try to alleviate the memory bottleneck in the LUT construction and distance calculation stages. In this subsection, we look at the performance issue in the top-k selection stage to further enhance MemANNS efficiency.
@@
Considered four threads, each maintaining a priority queue of six items. A max heap is employed to construct the priority queues. Once all four threads have completed their maintenance of the max heaps (Barrier 3), they can use semaphores to concurrently insert values from their thread‐local heaps into the total top‐k. To minimize unnecessary insertion, we propose to convert the thread-local max heaps into min heaps. We use the sem_take() and sem_give() semaphore functions to manage concurrent insertions of the top element from each min heap into the total result max heap. For any thread‐local min heap, if the top value is greater than the maximum value in the total max heap, it indicates that the remaining values in the thread‐local heap cannot contribute to the overall top‐k and can therefore be pruned.
@@ 5. Evaluation
5.1. Experimental Setup
Compared Baselines. We compare PIM-based MemANNS with CPU-based and GPU-based IVFPQ methods implemented using Faiss (Facebook AI Research, [n. d.]), a popular ANNS library developed by Meta. There are other advanced works to accelerate ANNS methods. FANNS (Jiang et al., 2023) employs FPGA to accelerate the IVFPQ method but is limited by its device memory, making it unsuitable for billion-scale searches. ANNA (Lee et al., 2022) designs specialized architecture to accelerate the IVFPQ method based on a simulator. Since MemANNS is based on real-world hardware, ANNA is not a suitable comparison. Juno (Liu et al., 2024) uses RT cores in GPUs to accelerate the IVFPQ method, but it achieves similar performance to Faiss on GPUs without RT cores, such as A100.
@@ 5.1. Experimental Setup
Hardware Setup. Table 2 shows the hardware specifications of the three compared solutions. For the CPU-based platform, we use two Intel Xeon Silver 4110@2.10GHz CPUs with 4 DDR4-2666Hz DRAM modules, providing 128 GB memory capacity and 85.3 GB/s bandwidth. We use one NVIDIA A100 GPU (Nvidia, 2020) with 80GB memory capacity and 1,935 GB/s bandwidth. For MemANNS, we use 7 UPMEME PIM modules, with 56 GB total memory capacity and 612.5 GB/s aggregated bandwidth. Existing study (Falevoz and Legriel, 2023) indicates that the peak power of each PIM DIMM is 23.22W, leading to 162W total peak power.
Benchmark. We evaluate the compared baselines using two billion-scale datasets, namely the SIFT1B (Jégou et al., 2011) and SPACEV1B (spa, 2021). The SIFT1B dataset is composed of 1 billion 128-dimensional vectors encoded into 16 dimensions. The SPACEV1B dataset contains 1 billion 100-dimensional vectors encoded into 20 dimensions. Both datasets contain 10,000 query vectors and corresponding ground truths for recall calculation. We process 1,000 queries at a time. The optimizations in MemANNS do not impact the recall.
@@ 5.2. Overall Performance Results
(a) QPS of CPU vs. MemANNS
(b) QPS of GPU vs MemANNS
(c) QPS/W of GPU vs. MemANNS
Figure 13. Normalized QPS and QPS/W of different ANNS solutions. $nprobe$ varies from 64, 128 to 256. #clusters varies from 4096, 8192 to 16384.
@@
We first compare the performance and efficiency of the three compared solutions. Results are shown in Figure 13. The term IVF4096 indicates that 1 billion points are partitioned into 4096 clusters. We vary nprobe from 64, 128 to 256, which indicates the number of clusters selected for each query search. All results are normalized to the values obtained from Faiss-CPU or Faiss-GPU when nprobe is set to 256 for each individual IVF.
@@
Compared to Faiss-CPU, MemANNS consistently achieves the highest queries per second (QPS) across all settings. Specifically, MemANNS accelerates query search times by 2.3x-4.3x for SIFT1B and by 2.1x-4.0x for SPACEV1B when compared to Faiss-CPU. Notably, under the same IVF setting, both MemANNS and Faiss-CPU experience a decrease in QPS as the nprobe increases, due to the increased search workload. Under the same nprobe, MemANNS obtains a higher QPS improvement compared to Faiss-CPU when the IVF increases. The reason is that a higher number of clusters results in smaller cluster sizes, which leads to fewer encoded points to search within each cluster and reduces data locality. Hence, the CPU, which has multiple cache layers, does not exhibit a linear increase in QPS with increasing IVF. In contrast, the DPU, which has a small-sized WRAM for cache and MRAM for main memory, is less affected by data locality, enabling MemANNS to achieve greater speedup as IVF increases.
@@
Comparing MemANNS with Faiss-GPU, Figure 13(b) shows the performance results in terms of QPS, and Figure 13(c) shows the energy efficiency results in terms of QPS per watt. Overall, MemANNS obtains comparable QPS as Faiss-GPU in most cases, except when IVF is 16384 and nprobe is 64. We used Nvidia Nsight (nsy, 2024) to look into the results and found that the significantly high performance of Faiss-GPU under this setting is because of the high parallelism of the top-k selection. The parallelism of this stage on SIFT1B is 9x higher than that on SPACEV1B under the same IVF and nprobe parameters. Note that the comparable performance of MemANNS and Faiss-GPU are obtained under a huge gap in the memory bandwidth and computing capabilities between PIM and A100 hardware, as detailed in Table 2.
@@
From the energy efficiency perspective, the 7 UPMEM DIMMs consume 162W peak power while A100 GPU consumes a significant 300W. Although the actual energy consumption during runtime differs from the peak power, we can use it as an approximation to compare the energy efficiency of MemANNS and Faiss-GPU. Results in Figure 13(c) show that, MemANNS achieves 2x higher QPS/W compared to Faiss-GPU in most cases, demonstrating better power efficiency. Beyond energy consumption, we also find that the per dollar QPS of MemANNS can be up to 9.3x higher than that of Faiss-GPU, indicating that MemANNS can significantly reduce costs in real production environments.
@@ 5.3. Scalability Study
Figure 14. Scalability of MemANNS under different #DPUs. Red triangles are measured from real hardware and red dash line represents predicted QPS. The yellow line indicates the QPS of Faiss-GPU. The blue vertical dash line shows when DPUs consume power equal to that of an A100 GPU.
Figure 15. The impact of MRAM read size.
@@
Existing systems can hold up to 20 UPMEM DIMMs, each containing 128 DPUs (Gómez-Luna et al., 2022), totaling 2560 DPUs. To study the performance of MemANNS beyond the seven PIM DIMMs we have, we evaluate the QPS of MemANNS with 500, 600, …, 900 DPUs and use the regression method to predict the QPS of MemANNS with up to 2560 DPUs. Due to the limited memory capacity when the number of DPUs is low, we evaluate the performance using the 500 million scale of the SIFT and SPACE datasets.
@@
As shown in Figure 15, the regression curve fits perfectly with the QPS results measured from 500-900 DPUs. When the number of DPUs increases, the QPS of MemANNS increases almost linearly, demonstrating good scalability of our system. When the number of DPUs is 2560, MemANNS can achieve up to 2.7x higher QPS compared to Faiss-GPU. Note that, even with 20 DIMMs, the cost of PIM is still much lower than that of A100 GPU (\$8000 vs. \$20,000). If we compare the performance of MemANNS and Faiss-GPU under the same peak power (i.e., 300W), we can use 1654 DPUs. As shown by the blue vertical dash line in Figure 15, MemANNS obtains higher QPS than Faiss-GPU at the same peak power constraint under all settings. This again demonstrates the energy efficiency of MemANNS, making it a practical solution for real-world large-scale production systems.
5.4. Sensitivity Study
In this subsection, we study the impact of several parameters to the performance of MemANNS, including the data size of each MRAM read, the number of threads during parallel execution in a single DPU and the required k size in the top-k query.
5.4.1. MRAM Read Size
As described in Section 4.2, the MRAM read latency does not increase linearly with size. To decide an optimal MRAM read size for each thread, we evaluate the performance of MemANNS with varied MRAM read sizes. Since the MRAM read size must be between 8 Bytes - 2048 Bytes and aligned with 8 Bytes, we varied the number of vectors read at once. For SIFT1B, we vary the number of vectors fetched in one MRAM read from 2, 4, …, to 64, making the MRAM read size vary from 64 Bytes, 128 Bytes, …, to 2 KB. For SPACEV1B, since the dimension of each encoded vector is 20, we vary the number of vectors in one MRAM read from 2, 4, …, to 32 to fit in the 2048 Bytes limit. We set IVF to 4096 and top-k to 10, testing on 14 threads. Figure 15 shows the evaluation results.
For both datasets, the QPS increases more rapidly when the number of vectors increases from 2 to 16, while becoming much more stable when the number of vectors exceeds 16. This is consistent with our observation in Figure 9, which shows that the MRAM read latency grows slowly when the data transfer size increases from 8 Bytes to 512 Bytes and grows dramatically beyond 512 Bytes. By default, we set the MRAM read size to 16 vectors to have good QPS and reasonable WRAM size at the same time.
5.4.2. Number of Threads per DPU
Figure 16. QPS of MemANNS as the #threads increases.
We evaluate the performance of MemANNS by varying the number of tasklets (namely threads in one DPU) from 1 to 24. Figure 16 shows our results for both datasets with different nprobe values. We normalized the values to those when the number of tasklets is one. We have similar observations in all settings. That is, the QPS increases linearly as the number of tasklets increases up to 11. Beyond 11 tasklets, the performance nearly saturates. The QPS of MemANNS with 11 tasklets is almost 11x higher than that with a single tasklet. This is because each DPU has a 14-stage pipeline, and only the last three stages can execute in parallel with the first two stages of the next instruction within the same thread. Using more than 11 tasklets makes full use of the pipeline and keeps the DPU busy. Thus, by default, we set the number of threads to 11 per DPU.
5.4.3. Top-K Size
(a) MemANNS vs. CPU
(b) MemANNS vs. GPU
Figure 17. The impact of top-k size
(a) SIFT1B
(b) SPACEV1B
Figure 18. Query search time breakdown
To study the impact of top-k size, we vary the k size from 1, 10 to 100 while keeping other parameters as default. We compare the QPS of MemANNS with that of Faiss-CPU, and the QPS/W of MemANNS with that of Faiss-GPU. Figure 17 shows the results. All results are normalized to the corresponding results of the top-100. We have the following observations.
First, given the same k size, MemANNS outperforms Faiss-CPU by 2.6x speedup in QPS on average. In terms of energy efficiency, MemANNS achieves more than 2x better QPS/W compared to Faiss-GPU. Second, under different k sizes, the QPS of Faiss-CPU remains almost unchanged while the QPS of MemANNS decreases slightly when k varies from 1 to 100. This is mainly due to the increased size of the top-k list, which causes higher CPU-DPU communication overhead. Faiss-GPU shares a similar trend in QPS change as MemANNS. This is because selecting top-k occupies the largest portion of query search time in Faiss-GPU due to the large amount of CUDA stream synchronization overhead.
We further breakdown the query search time of the compared solutions under different k sizes, as shown in Figure 18. We have the following observations. First, comparing different hardware under the same k size, MemANNS reduces the ratio of the distance calculation stage from 99.5% on Faiss-CPU to 75.5%, indicating that MemANNS is effectively alleviating the memory bottleneck in the CPU-based implementation. Although the GPU is empowered with a much larger memory bandwidth compared to 7 PIM DIMMs, helping to address memory access issues, it incurs considerable CUDA stream synchronization while identifying top-k. This synchronization has emerged as a new bottleneck, consuming more than 85% of processing time for large-scale datasets. In contrast, DPUs do not share this bottleneck and demonstrate performance comparable to that of the GPUs. Secondly, comparing different k sizes, the time spent on the top-k selection stage increases from 84% to 89% on GPU and from 9% to 17% on DPU. The ratio of this stage on CPU remains unchanged due to the major time taken by distance calculation. This is consistent with our analysis of the impact of top-k sizes discussed above.
5.5. Summary and Discussion
The experimental results underscore the advantages of MemANNS in terms of performance, energy efficiency and scalability. Its ability to outperform Faiss-CPU and maintain competitive performance with Faiss-GPU, coupled with its superior energy and cost efficiency, makes MemANNS a compelling solution for large-scale similarity search applications. However, UPMEM reports a maximum integer processing capacity of 896 GOPS (Gómez-Luna et al., 2022), whereas GPUs like A100 offer significantly higher computational power at 156 TFLOPS (Nvidia, 2020). In addition, it is essential to mitigate the huge latency gap between multiplication and addition operations in DPU. As UPMEM is the first commercial PIM hardware to enter the market, future versions are expected to enhance computational capacity.

### v2

@@ 4.4. Top-K Pruning
The above optimizations try to alleviate the memory bottleneck in the LUT construction and distance calculation stages. In this subsection, we look at the performance issue in the top-k selection stage to further enhance UpANNS efficiency.
@@
Considered four threads, each maintaining a priority queue of six elements. A max heap is employed to construct the priority queues. Once all four threads have completed their maintenance of the max heaps (Barrier 3), they can use semaphores to concurrently insert values from their thread‐local heaps into the total top‐k. To minimize unnecessary insertion, we propose to convert the thread-local max heaps into min heaps. We use the sem_take() and sem_give() semaphore functions to manage concurrent insertions of the top element from each min heap into the total result max heap. For any thread‐local min heap, if the top value is greater than the maximum value in the total max heap, it indicates that the remaining values in the thread‐local heap cannot contribute to the overall top‐k and can therefore be pruned.
@@ 5. Evaluation
5.1. Experimental Setup
Compared Baselines. We compare UpANNS with three baselines: PIM-naive, Faiss-CPU and Faiss-GPU. PIM-naive is the naive implementation of IVFPQ on PIM with our PIM resource management strategy. Faiss-CPU and Faiss-GPU adopt the CPU- and GPU-based implementations of IVFPQ from Faiss library (Facebook AI Research, [n. d.]), the most popular ANNS library developed by Meta.
While there are other advanced works to accelerate ANNS methods, their methodologies or hardware compatibility differ fundamentally from UpANNS. For example, FANNS (Jiang et al., 2023a) accelerates IVFPQ using FPGAs but faces device memory constraints, limiting scalability to billion-scale datasets. Some studies (Lee et al., 2022a) optimizes IVFPQ via simulator-based architectures, thus are incomparable to UpANNS, which targets real-world hardware. To the best of our knowledge, no accessible PIM-based ANNS implementation exists for direct comparison. For example, Proxima (Xu et al., 2023) estimated its performance using an in-house simulator. To bridge this gap, we implemented a rigorous naive PIM baseline (PIM-naive) that serves as a foundation for UpANNS. Notably, the above works (Lee et al., 2022a; Jiang et al., 2023a; Xu et al., 2023) all use Faiss as the state-of-the-art method, making it a reasonable choice for comparison.
Figure 10. QPS of different ANNS solutions normalized to that of Faiss-CPU when IVF is 4096 and nprobe is 256. nprobe varies from 64, 128 to 256. #clusters varies from 4096, 8192 to 16384.
Figure 11. The ratio of maximum process and average process improvement due to the PIM-Aware Workload Distribution strategy under different nprobe and IVF settings.
(a) QPS of GPU vs UpANNS
(b) QPS/W of GPU vs. UpANNS
Figure 12. QPS and QPS/W of Faiss-GPU and UpANNS, normalized to those of Faiss-GPU when IVF is 4096 and nprobe is 256. nprobe varies from 64, 128 to 256. #clusters varies from 4096, 8192 to 16384. The results are normalized to in SIFT1B and SPACEV1B, nprobe is 64 in DEEP1B.
Hardware Setup. Table 1 shows the hardware specifications of the four compared solutions. Except for the components noted, the remaining hardware is the same across all setups. For the CPU-based platform, we use two Intel Xeon Silver 4110@2.10GHz CPUs with 4 DDR4-2666Hz DRAM modules, providing 128 GB memory capacity and 85.3 GB/s bandwidth. We use one NVIDIA A100 GPU (Nvidia, 2020) with 80GB memory capacity and 1,935 GB/s bandwidth. For PIM-naive and UpANNS, we use 7 UPMEME PIM modules, with 56 GB total memory capacity and 612.5 GB/s aggregated bandwidth. Existing study (Falevoz and Legriel, 2023) indicates that the peak power of each PIM DIMM is 23.22W, leading to 162W total peak power.
Benchmark. We evaluate the compared baselines using three billion-scale datasets, namely the DEEP1B (Babenko and Lempitsky, 2016), SIFT1B (Jégou et al., 2011) and SPACEV1B (spa, 2021). The DEEP1B dataset contains 1 billion 96-dimensional vectors encoded into 12 dimensions. The SIFT1B dataset contains 1 billion 128-dimensional vectors encoded into 16 dimensions. The SPACEV1B dataset contains 1 billion 100-dimensional vectors encoded into 20 dimensions. We process 1,000 queries at a time. The optimizations in UpANNS do not impact the accuracy.
@@ 5.2. Overall Performance Results
We first compare the performance and efficiency of the four compared solutions. Results are shown in Figure 11 and Figure 12. The term IVF4096 indicates that 1 billion points are partitioned into 4096 clusters. We vary nprobe from 64, 128 to 256, which indicates the number of clusters selected for each query search. All results are normalized to the values obtained from Faiss-CPU or Faiss-GPU when nprobe=256 and IVF=4096 for each dataset. For DEEP1B, due to GPU out-of-memory errors (marked with blue ’X’ in Figure 12), results are normalized to Faiss-GPU with nprobe=64 and IVF=4096.
@@
Compared to Faiss-CPU and PIM-naive, UpANNS consistently achieves the highest queries per second (QPS) across all settings. Specifically, UpANNS accelerates query search times by 1.6x-3.8x for DEEP1B, by 2.3x-4.3x for SIFT1B and by 2.1x-4.0x for SPACEV1B when compared to Faiss-CPU. Notably, under the same IVF setting, all solutions experience a decrease in QPS as nprobe increases, due to the increased search workload. Under the same nprobe, UpANNS obtains a higher QPS improvement compared to Faiss-CPU when the IVF increases. The reason is that a higher number of clusters results in smaller cluster sizes, which leads to fewer encoded points to search within each cluster and reduces data locality. Hence, the CPU, which has multiple cache layers, does not exhibit a linear increase in QPS with increasing IVF. In contrast, the DPU, which has a small-sized WRAM for cache and MRAM for main memory, is less affected by data locality, enabling UpANNS to achieve greater speedup as IVF increases. Although PIM-naive also surpasses Faiss-CPU, it underperforms UpANNS by up to 3.1x, underscoring the critical role of architectural optimizations in the proposed solution.
@@
Comparing UpANNS with Faiss-GPU, Figure 12(a) shows the performance results in terms of QPS, and Figure 12(b) shows the energy efficiency results in terms of QPS per watt. Overall, UpANNS obtains comparable QPS as Faiss-GPU in most cases, except when IVF is 16384 and nprobe is 64. We used Nvidia Nsight (nsy, 2024) to look into the results and found that the significantly high performance of Faiss-GPU under this setting is because of the different parallelism of the top-k selection. The parallelism of this stage on SIFT1B is 9x higher than the one on SPACEV1B under the same IVF and nprobe parameters. Note that the comparable performance of UpANNS and Faiss-GPU are obtained under a huge gap in the memory bandwidth and computing capabilities between PIM and A100 hardware, as detailed in Table 1.
@@
From the energy efficiency perspective, the 7 UPMEM DIMMs consume 162W peak power while A100 GPU consumes a significant 300W. Although the actual energy consumption during runtime differs from the peak power, we can use it as an approximation to compare the energy efficiency of UpANNS and Faiss-GPU. Results in Figure 12(b) show that, UpANNS achieves 2x higher QPS/W compared to Faiss-GPU in most cases, demonstrating better power efficiency. While newer GPU models like H100 SXM ( 3.5TB/s bandwidth, 700W) and GH200 ( 4.9TB/s bandwidth, 1000W) provide 2-3× higher bandwidth, their power requirements scale proportionally, reinforcing UpANNS’s position as the more energy-efficient solution. Beyond energy consumption, we also find that the per dollar QPS of UpANNS can be up to 9.3x higher than that of Faiss-GPU, indicating that UpANNS can significantly reduce costs in real production environments.
5.3. Ablation Study
In this subsection, we conduct ablation studies to evaluate the effectiveness of our key optimizations in UpANNS, including the PIM-Aware Workload Distribution, multi-threading design, Co-occurrence Aware Encoding, and Top-K Pruning strategy.
5.3.1. PIM-Aware Workload Distribution
Since the workloads need to be distributed among DPUs and executed in parallel, the largest workload among DPUs determines the overall performance. We evaluate the performance improvement due to the PIM-Aware Workload Distribution strategy compared to the naive distribution strategy that assigns clusters randomly to DPUs. Figure 11 shows the ratio of maximum process to average process under different nprobe and IVF settings. A ratio closer to 1 indicates that the workload is more evenly distributed among DPUs. We observe that the ratio for PIM-naive is significantly higher than 1, especially when the IVF and nprobe are small. In contrast, the ratio for UpANNS is close to 1 under all settings for all datasets, indicating that the PIM-Aware Workload Distribution strategy effectively balances the workload among DPUs.
5.3.2. Number of Threads per DPU
Figure 13. QPS of UpANNS as the #threads increases
UPMEM PIM could support up to 24 threads per DPU (upm, 2023a). Therefore, we evaluate the performance of UpANNS by varying the number of threads from 1 to 24. Figure 13 shows our results for three datasets with different nprobe values. We normalize the values to those when the number of tasklets is one. We have similar observations in all settings. That is, the QPS increases linearly as the number of tasklets increases up to 11. Beyond 11 tasklets, the performance nearly saturates. The QPS of UpANNS with 11 tasklets is almost 11x higher than that with a single tasklet. This is because each DPU has a 14-stage pipeline, and only the last three stages can execute in parallel with the first two stages of the next instruction within the same thread. Using more than 11 tasklets makes full use of the pipeline and keeps the DPU busy. Thus, by default, we set #threads to 11 per DPU.
5.3.3. Co-occurrence Aware Encoding
Figure 14. Performance improvement due to Co-occurrence Aware Encode strategy (CAE) under different nprobe and corresponding length reduction rates.
To evaluate the performance improvement due to the Co-occurrence Aware Encoding strategy, we select queries that choose top-nprobe clusters with the highest vector length reduction rates and calculate the average vector length reduction rate in the maximum workload DPU. Figure 14 shows the performance improvement due to the Co-occurrence Aware Encoding strategy under different nprobe values and corresponding length reduction rates. We observe the following for all settings. First, the performance improvement correlates positively with the length reduction rate, which is consistent with our design. Second, the LUT construction time increases slightly due to the additional time needed to construct the partial sum. Third, the performance improvement becomes more significant with higher length reduction rates. This is because a higher length reduction rate means less data needs to be transferred between MRAM and WRAM, allowing calculations to be done more efficiently. The time breakdown also shows that distance calculation time decreases more when the length reduction rate is higher.
5.3.4. Top-K Pruning
Figure 15. Time reduction with proposed top-k selection strategy.
Figure 16. The impact of different batch size on query latency.
We evaluate the time of the top-k selection stage with and without the proposed top-k selection strategy. Figure 16 shows the time reduction due to our proposed strategy. The results are normalized to the time with the proposed strategy in top-10. We observe that the top-k selection time increases linearly with the top-k value. The proposed strategy significantly reduces the time for top-k selection, especially when the top-k value is large. This is because our strategy reduces the number of comparisons needed to find the top-k results.
5.4. Sensitivity Study
We study the impact of three parameters to the performance of UpANNS, including the batch size of queries, the data size of each MRAM read and the required k size in the top-k query.
5.4.1. Batch Size
Figure 16 shows the impact of different batch sizes on query latency. We set the IVF to 4096 and nprobe to 64, varying the batch size (#BS) from 10, 100 to 1000. We observe that UpANNS achieves the lowest query latency in all settings, and the speedup of UpANNS over Faiss-CPU and PIM-naive increases as the batch size increases. This is because the overhead of the pre-processing and post-processing stages is amortized over a larger number of queries. This makes it suitable for real-world applications where the number of queries is usually large.
5.4.2. MRAM Read Size
As described in Section 4.2, the MRAM read latency does not increase linearly with size. To decide an optimal MRAM read size for each thread, we evaluate the performance of UpANNS with varied MRAM read sizes. Since the MRAM read size must be between 8 Bytes - 2048 Bytes and aligned with 8 Bytes, we varied the number of vectors read at once. For example, we vary the number of vectors fetched in one MRAM read from 2, 4, …, to 64, making the MRAM read size vary from 64 Bytes, 128 Bytes, …, to 2 KB to fit in the 2048 Bytes limit in SIFT1B. We set IVF to 4096 and top-k to 10. Figure 18 shows the evaluation results.
For all datasets, the QPS increases more rapidly when the number of vectors increases from 2 to 16, while becoming much more stable when the number of vectors exceeds 16. This is consistent with our observation in Figure 7, which shows that the MRAM read latency grows slowly when the data transfer size increases from 8 Bytes to 512 Bytes and grows dramatically beyond 512 Bytes. By default, we set the MRAM read size to 16 vectors to have good QPS and reasonable WRAM size at the same time.
5.4.3. Top-K Size
Figure 17. The impact of MRAM read size.
Figure 18. The impact of top-k size.
Figure 19. Query search time breakdown
To evaluate the impact of top-k size, we vary k from 1 to 100 while maintaining default parameters. Figure 18 presents normalized QPS results (relative to Faiss-CPU’s top-100 performance) for UpANNS, Faiss-CPU, and Faiss-GPU. UpANNS achieves consistent performance advantages, achieving an average 2.5x higher QPS than Faiss-CPU and 1.6x higher than Faiss-GPU across datasets. While Faiss-CPU’s QPS remains stable across k values, UpANNS and Faiss-GPU exhibit slight QPS degradation as k increases. This stems from the growing top-k list size, which amplifies CPU-DPU communication costs for UpANNS and CUDA stream synchronization overhead for Faiss-GPU during top-k selection.
A detailed processing time breakdown (Figure 19) reveals critical insights. UpANNS significantly reduces the distance calculation stage’s time ratio from 99.5% (Faiss-CPU) to 75.5–80% across datasets, demonstrating its effectiveness in mitigating CPU memory bottlenecks. While GPUs leverage high memory bandwidth to alleviate memory access challenges, their performance is constrained by CUDA synchronization during identifying top-k, consuming over 85% of processing time for large datasets. DPUs avoid this bottleneck, achieving GPU-competitive performance. As k increases, the top-k stage’s time ratio grows from 76% to 89% on GPUs and 9% to 17% on DPUs, while remaining negligible on CPUs due to their dominance by distance calculations. These trends align with the observed QPS variations under different k sizes.
@@ 5.5. Scalability Study
[…]
This figure shows the normalized QPS of different solutions with varying nprobe and cluster counts. Refer to caption
@@
Figure 20. Scalability of UpANNS under different #DPUs. Red triangles are measured from real hardware and red dash line represents predicted QPS. The yellow line indicates the QPS of Faiss-GPU. The blue vertical dash line shows when DPUs consume power equal to that of an A100 GPU.
@@
Existing systems can hold up to 20 UPMEM DIMMs, each containing 128 DPUs (Gómez-Luna et al., 2022), totaling 2560 DPUs. To study the performance of UpANNS beyond the seven PIM DIMMs we have, we evaluate the QPS of UpANNS with 500, 600, …, 900 DPUs and use the regression method to predict the QPS of UpANNS with up to 2560 DPUs. Due to the limited memory capacity when the number of DPUs is low, we evaluate the performance using 500 million scale.
@@
As shown in Figure 20, the regression curve fits perfectly with the QPS results measured from 500-900 DPUs. When the number of DPUs increases, the QPS of UpANNS increases almost linearly, demonstrating good scalability of our system. When the number of DPUs is 2560, UpANNS can achieve up to 2.6x higher QPS compared to Faiss-GPU. Note that, even with 20 DIMMs, the cost of PIM is still much lower than that of A100 GPU (\$8000 vs. \$20,000). If we compare the performance of UpANNS and Faiss-GPU under the same peak power (i.e., 300W), we can use 1654 DPUs. As shown by the blue vertical dash line in Figure 20, UpANNS obtains higher QPS than Faiss-GPU at the same peak power constraint under all settings. This again demonstrates the energy efficiency of UpANNS, making it practical for real-world large-scale production systems. While our current evaluation focuses on single-host configurations, UpANNS can be easily extended to multi-host configurations. Only query distribution and result aggregation require cross-host communication. The core memory-intensive search operations remain local to each host, ensuring efficient scalability.